OpenStudy (anonymous):

Mulitiply and Simplify by factoring 26^3 and 14c^14

OpenStudy (anonymous):



OpenStudy (campbell_st):

\[\sqrt{26^3} \times \sqrt{14c^4} = \sqrt{26^3 \times 14c^4}\] will give \[\sqrt{ 26^2 \times c^4 \times26 \times 14}\] \[26c^2 \sqrt{364}\] \[26c^2 \sqrt{364} = 26c^2 \times 2 \sqrt{91}\] \[52c^2\sqrt{91}\]

OpenStudy (anonymous):

so the final answer to plug in is the last one right?

OpenStudy (anonymous):

your answer was wrong, it says the correct answer is 2bc^2 sqrt 7b